# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN)

### v1

@@ Acknowledgements
The breakdown of systematic uncertainty contributions for the reported observables is shown in Table End Matter.
@@
Stat. Syst. Background Efficiency Background Fixed model Spin-2
(total) fractions map modeling parameters components
$\mathcal{A}_{{C\!P}}$ 0.062 0.017 0.0001 0.0049 0.0066 0.0127 0.0067
$\mathcal{A}_{{C\!P}}(A_{0})$ 0.083 0.025 0.0003 0.0061 0.0114 0.0152 0.0151
$\mathcal{A}_{{C\!P}}(A_{\parallel})$ 0.137 0.024 0.0015 0.0040 0.0175 0.0107 0.0116
$\mathcal{A}_{{C\!P}}(A_{\perp})$ 0.140 0.022 0.0005 0.0035 0.0158 0.0086 0.0116
$\Delta_{{C\!P}}(A_{0})$ 0.177 0.048 0.0031 0.0129 0.0304 0.0320 0.0141
$\Delta_{{C\!P}}(A_{\parallel})$ 0.187 0.040 0.0040 0.0021 0.0138 0.0247 0.0282
$\Delta_{{C\!P}}(A_{\perp})$ 0.180 0.048 0.0013 0.0053 0.0175 0.0257 0.0365
Results of the fitted amplitude parameters, i.e. the real and imaginary parts of amplitude couplings, are summarized in Table 2. Parameters marked with “—” are constrained in the fit, as explained below.
@@
$\mathcal{R}e(A_{\rho({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $0.9735$ $\pm 0.0314$ $\pm 0.0121$
@@
$\mathcal{R}e(A_{\rho{{K}^{*}}}^{0})$ $0.1641$ $\pm 0.0254$ $\pm 0.0453$ $0.2392$ $\pm 0.0332$ $\pm 0.0749$
$\mathcal{I}m(A_{\rho{{K}^{*}}}^{0})$ $0.0242$ $\pm 0.0266$ $\pm 0.0350$ $0.2811$ $\pm 0.0326$ $\pm 0.0773$
$\mathcal{R}e(A_{\rho{{K}^{*}}}^{\parallel})$ 0.1289-0.1289 $\pm 0.0141$ $\pm 0.0329$ 0.1203-0.1203 $\pm 0.0143$ $\pm 0.0294$
$\mathcal{I}m(A_{\rho{{K}^{*}}}^{\parallel})$ $0.0297$ $\pm 0.0184$ $\pm 0.0258$ 0.0308-0.0308 $\pm 0.0180$ $\pm 0.0230$
$\mathcal{R}e(A_{\rho{{K}^{*}}}^{\perp})$ 0.1021-0.1021 $\pm 0.0142$ $\pm 0.0292$ 0.1385-0.1385 $\pm 0.0133$ $\pm 0.0331$
$\mathcal{I}m(A_{\rho{{K}^{*}}}^{\perp})$ $0.0246$ $\pm 0.0160$ $\pm 0.0211$ 0.0246-0.0246 $\pm 0.0174$ $\pm 0.0269$
$\mathcal{R}e(A_{\omega{{K}^{*}}}^{0})$ $0.0005$ $\pm 0.0067$ $\pm 0.0024$ 0.0087-0.0087 $\pm 0.0056$ $\pm 0.0032$
$\mathcal{I}m(A_{\omega{{K}^{*}}}^{0})$ 0.0114-0.0114 $\pm 0.0058$ $\pm 0.0030$ 0.0119-0.0119 $\pm 0.0074$ $\pm 0.0028$
$\mathcal{R}e(A_{\omega{{K}^{*}}}^{\parallel})$ 0.0023-0.0023 $\pm 0.0030$ $\pm 0.0009$ 0.0006-0.0006 $\pm 0.0023$ $\pm 0.0006$
$\mathcal{I}m(A_{\omega{{K}^{*}}}^{\parallel})$ 0.0025-0.0025 $\pm 0.0028$ $\pm 0.0007$ $0.0020$ $\pm 0.0022$ $\pm 0.0006$
$\mathcal{R}e(A_{\omega{{K}^{*}}}^{\perp})$ 0.0001-0.0001 $\pm 0.0044$ $\pm 0.0008$ 0.0027-0.0027 $\pm 0.0033$ $\pm 0.0007$
$\mathcal{I}m(A_{\omega{{K}^{*}}}^{\perp})$ 0.0042-0.0042 $\pm 0.0030$ $\pm 0.0013$ $0.0015$ $\pm 0.0034$ $\pm 0.0006$
$\mathcal{R}e(A_{\omega({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $0.0052$ $\pm 0.0112$ $\pm 0.0044$ 0.0001-0.0001 $\pm 0.0119$ $\pm 0.0018$
$\mathcal{I}m(A_{\omega({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $0.0254$ $\pm 0.0096$ $\pm 0.0022$ $0.0276$ $\pm 0.0087$ $\pm 0.0024$
$\mathcal{R}e(A_{f_{0}(500){{K}^{*}}})$ $1.5539$ $\pm 0.1852$ $\pm 0.6291$ $1.0127$ $\pm 0.2455$ $\pm 0.3860$
$\mathcal{I}m(A_{f_{0}(500){{K}^{*}}})$ $0.3152$ $\pm 0.3778$ $\pm 0.2455$ $1.1808$ $\pm 0.2386$ $\pm 0.3082$
$\mathcal{R}e(A_{f_{0}(980){{K}^{*}}})$ 0.1730-0.1730 $_{-0.1102}^{+0.0909}$ $_{-0.2839}^{+0.2468}$ 0.3988-0.3988 $_{-0.1624}^{+0.0667}$ $_{-0.4019}^{+0.4691}$
$\mathcal{I}m(A_{f_{0}(980){{K}^{*}}})$ 0.6598-0.6598 $_{-0.3013}^{+0.0811}$ $_{-0.3960}^{+0.4169}$ 0.5940-0.5940 $_{-0.2850}^{+0.0824}$ $_{-0.4409}^{+0.5030}$
$\mathcal{R}e(A_{f_{0}(1370){{K}^{*}}})$ $0.4861$ $_{-0.3195}^{+0.3206}$ $_{-0.3780}^{+0.5969}$ $1.0387$ $_{-0.2688}^{+0.1879}$ $_{-0.4077}^{+0.4365}$
$\mathcal{I}m(A_{f_{0}(1370){{K}^{*}}})$ $0.1711$ $_{-0.2424}^{+0.2860}$ $_{-0.5106}^{+0.6180}$ $0.4089$ $_{-0.2555}^{+0.2378}$ $_{-0.7276}^{+0.9128}$
$\mathcal{R}e(A_{f_{0}(500)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $0.6749$ $\pm 0.4192$ $\pm 0.3066$ $1.3057$ $\pm 0.3611$ $\pm 0.4932$
$\mathcal{I}m(A_{f_{0}(500)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ 0.3747-0.3747 $\pm 0.4989$ $\pm 0.3302$ 0.1689-0.1689 $\pm 0.5528$ $\pm 0.2579$
$\mathcal{R}e(A_{f_{0}(980)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ 0.2589-0.2589 $_{-0.2642}^{+0.0870}$ $_{-0.3531}^{+0.4903}$ 0.3503-0.3503 $_{-0.2573}^{+0.0924}$ $_{-0.1454}^{+0.4844}$
$\mathcal{I}m(A_{f_{0}(980)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $1.3425$ $_{-0.2222}^{+0.6411}$ $_{-0.3616}^{+0.3440}$ $0.9277$ $_{-0.1349}^{+0.4909}$ $_{-0.3160}^{+0.3220}$
$\mathcal{R}e(A_{f_{0}(1370)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $1.3519$ $_{-0.5138}^{+0.5718}$ $_{-0.7381}^{+1.2252}$ $0.2608$ $_{-0.5300}^{+0.5436}$ $_{-0.7116}^{+0.6526}$
$\mathcal{I}m(A_{f_{0}(1370)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $0.0270$ $_{-0.4300}^{+1.0180}$ $_{-0.9053}^{+0.8738}$ $0.6014$ $_{-0.4542}^{+0.4970}$ $_{-0.5779}^{+0.4421}$
$\mathcal{R}e(A_{a_{1}(1260)(\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{0})$ 2.5802-2.5802 $\pm 1.3487$ $\pm 1.2300$ 2.0785-2.0785 $\pm 0.6787$ $\pm 0.9465$
$\mathcal{I}m(A_{a_{1}(1260)(\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{0})$ 0.5676-0.5676 $\pm 0.9027$ $\pm 0.6581$ 0.8797-0.8797 $\pm 0.6419$ $\pm 0.6604$
$\mathcal{R}e(A_{a_{1}(1260)(\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{\parallel})$ 0.1507-0.1507 $\pm 1.5943$ $\pm 0.8349$
$\mathcal{I}m(A_{a_{1}(1260)(\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{\parallel})$ $3.0275$ $\pm 1.3924$ $\pm 1.0201$
$\mathcal{R}e(A_{a_{1}(1260)(\rightarrow f_{0}(500)\pi){{K}^{0}_{\mathrm{S}}}})$ 8.6330-8.6330 $_{-1.8408}^{+1.8330}$ $_{-2.8059}^{+9.6604}$
$\mathcal{I}m(A_{a_{1}(1260)(\rightarrow f_{0}(500)\pi){{K}^{0}_{\mathrm{S}}}})$ $3.4232$ $_{-0.7814}^{+1.2376}$ $_{-0.6673}^{+0.9525}$
$\mathcal{R}e(A_{a_{1}(1260)(\rightarrow f_{0}(980)\pi){{K}^{0}_{\mathrm{S}}}})$ $1.2550$ $_{-1.2627}^{+0.8240}$ $_{-1.2970}^{+1.4950}$
$\mathcal{I}m(A_{a_{1}(1260)(\rightarrow f_{0}(980)\pi){{K}^{0}_{\mathrm{S}}}})$ $0.7388$ $_{-1.0314}^{+0.8666}$ $_{-0.3494}^{+0.4398}$
$\mathcal{R}e(A_{a_{1}(1640)(\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{0})$ 0.3555-0.3555 $\pm 0.4472$ $\pm 0.1914$ 0.2564-0.2564 $\pm 0.4432$ $\pm 0.1489$
$\mathcal{I}m(A_{a_{1}(1640)(\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{0})$ 0.0363-0.0363 $\pm 0.5007$ $\pm 0.2577$ 0.0859-0.0859 $\pm 0.3758$ $\pm 0.1804$
$\mathcal{R}e(A_{a_{1}(1640)(\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{\parallel})$ 0.3892-0.3892 $\pm 0.8504$ $\pm 0.1917$
$\mathcal{I}m(A_{a_{1}(1640)(\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{\parallel})$ 0.2174-0.2174 $\pm 0.9290$ $\pm 0.4535$
$\mathcal{R}e(A_{a_{1}(1640)(\rightarrow f_{0}(500)\pi){{K}^{0}_{\mathrm{S}}}})$ $3.8584$ $\pm 1.3145$ $\pm 0.7881$
$\mathcal{I}m(A_{a_{1}(1640)(\rightarrow f_{0}(500)\pi){{K}^{0}_{\mathrm{S}}}})$ $0.5558$ $\pm 1.1537$ $\pm 0.5703$
$\mathcal{R}e(A_{a_{1}(1640)(\rightarrow f_{0}(980)\pi){{K}^{0}_{\mathrm{S}}}})$ 0.3051-0.3051 $_{-0.7193}^{+0.6560}$ $_{-0.7802}^{+0.6627}$
$\mathcal{I}m(A_{a_{1}(1640)(\rightarrow f_{0}(980)\pi){{K}^{0}_{\mathrm{S}}}})$ 1.8068-1.8068 $_{-0.6458}^{+0.6287}$ $_{-0.9045}^{+0.7290}$

### v2

@@ Acknowledgements
The breakdown of systematic uncertainty contributions for the reported observables is shown in Table 1.
@@
Stat. Syst. Background Efficiency Background Mass Spin-2
(total) fractions map modeling models components
$\mathcal{A}_{{C\!P}}$ 0.062 0.024 0.0001 0.0049 0.0066 0.0216 0.0067
$\mathcal{A}_{{C\!P}}(A_{0})$ 0.083 0.029 0.0003 0.0061 0.0114 0.0206 0.0151
$\mathcal{A}_{{C\!P}}(A_{\parallel})$ 0.137 0.027 0.0015 0.0040 0.0175 0.0166 0.0116
$\mathcal{A}_{{C\!P}}(A_{\perp})$ 0.140 0.051 0.0005 0.0035 0.0158 0.0467 0.0116
$\Delta_{{C\!P}}(A_{0})$ 0.177 0.048 0.0031 0.0129 0.0304 0.0314 0.0141
$\Delta_{{C\!P}}(A_{\parallel})$ 0.187 0.114 0.0040 0.0021 0.0138 0.1092 0.0282
$\Delta_{{C\!P}}(A_{\perp})$ 0.180 0.122 0.0013 0.0053 0.0175 0.1145 0.0365
Results of the fitted amplitude parameters, i.e. the real and imaginary parts of amplitude couplings, are summarized in Table 2. Asymmetric uncertainties are evaluated following the procedure described in Ref. [79]. Parameters marked with “—” are constrained in the fit, as explained below.
@@
$\mathcal{R}e(A_{\rho({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $0.9735$ $\pm 0.0314$ $\pm 0.0142$
@@
$\mathcal{R}e(A_{\rho{{K}^{*}}}^{0})$ $0.1641$ $\pm 0.0254$ $\pm 0.0415$ $0.2392$ $\pm 0.0332$ $\pm 0.0750$
$\mathcal{I}m(A_{\rho{{K}^{*}}}^{0})$ $0.0242$ $\pm 0.0266$ $\pm 0.0316$ $0.2811$ $\pm 0.0326$ $\pm 0.0779$
$\mathcal{R}e(A_{\rho{{K}^{*}}}^{\parallel})$ 0.1289-0.1289 $\pm 0.0141$ $\pm 0.0332$ 0.1203-0.1203 $\pm 0.0143$ $\pm 0.0293$
$\mathcal{I}m(A_{\rho{{K}^{*}}}^{\parallel})$ $0.0297$ $\pm 0.0184$ $\pm 0.0307$ 0.0308-0.0308 $\pm 0.0180$ $\pm 0.0209$
$\mathcal{R}e(A_{\rho{{K}^{*}}}^{\perp})$ 0.1021-0.1021 $\pm 0.0142$ $\pm 0.0261$ 0.1385-0.1385 $\pm 0.0133$ $\pm 0.0325$
$\mathcal{I}m(A_{\rho{{K}^{*}}}^{\perp})$ $0.0246$ $\pm 0.0160$ $\pm 0.0249$ 0.0246-0.0246 $\pm 0.0174$ $\pm 0.0248$
$\mathcal{R}e(A_{\omega{{K}^{*}}}^{0})$ $0.0005$ $\pm 0.0067$ $\pm 0.0059$ 0.0087-0.0087 $\pm 0.0056$ $\pm 0.0034$
$\mathcal{I}m(A_{\omega{{K}^{*}}}^{0})$ 0.0114-0.0114 $\pm 0.0058$ $\pm 0.0126$ 0.0119-0.0119 $\pm 0.0074$ $\pm 0.0029$
$\mathcal{R}e(A_{\omega{{K}^{*}}}^{\parallel})$ 0.0023-0.0023 $\pm 0.0030$ $\pm 0.0009$ 0.0006-0.0006 $\pm 0.0023$ $\pm 0.0005$
$\mathcal{I}m(A_{\omega{{K}^{*}}}^{\parallel})$ 0.0025-0.0025 $\pm 0.0028$ $\pm 0.0013$ $0.0020$ $\pm 0.0022$ $\pm 0.0007$
$\mathcal{R}e(A_{\omega{{K}^{*}}}^{\perp})$ 0.0001-0.0001 $\pm 0.0044$ $\pm 0.0014$ 0.0027-0.0027 $\pm 0.0033$ $\pm 0.0008$
$\mathcal{I}m(A_{\omega{{K}^{*}}}^{\perp})$ 0.0042-0.0042 $\pm 0.0030$ $\pm 0.0013$ $0.0015$ $\pm 0.0034$ $\pm 0.0007$
$\mathcal{R}e(A_{\omega({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $0.0052$ $\pm 0.0112$ $\pm 0.0041$ 0.0001-0.0001 $\pm 0.0119$ $\pm 0.0019$
$\mathcal{I}m(A_{\omega({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $0.0254$ $\pm 0.0096$ $\pm 0.0048$ $0.0276$ $\pm 0.0087$ $\pm 0.0026$
$\mathcal{R}e(A_{f_{0}(500){{K}^{*}}})$ $1.5539$ $\pm 0.1852$ $\pm 0.3636$ $1.0127$ $\pm 0.2455$ $\pm 0.2885$
$\mathcal{I}m(A_{f_{0}(500){{K}^{*}}})$ $0.3152$ $\pm 0.3778$ $\pm 0.2541$ $1.1808$ $\pm 0.2386$ $\pm 0.3634$
$\mathcal{R}e(A_{f_{0}(980){{K}^{*}}})$ 0.1784-0.1784 $_{-0.1102}^{+0.0909}$ $^{+0.1183}_{-0.1341}$ 0.4046-0.4046 $_{-0.1624}^{+0.0667}$ $^{+0.0913}_{-0.1537}$
$\mathcal{I}m(A_{f_{0}(980){{K}^{*}}})$ 0.6713-0.6713 $_{-0.3013}^{+0.0811}$ $^{+0.1309}_{-0.1595}$ 0.6048-0.6048 $_{-0.2850}^{+0.0824}$ $^{+0.1284}_{-0.1519}$
$\mathcal{R}e(A_{f_{0}(1370){{K}^{*}}})$ $0.5053$ $_{-0.3195}^{+0.3206}$ $^{+0.2821}_{-0.2658}$ $1.0462$ $_{-0.2688}^{+0.1879}$ $^{+0.3300}_{-0.2286}$
$\mathcal{I}m(A_{f_{0}(1370){{K}^{*}}})$ $0.1557$ $_{-0.2424}^{+0.2860}$ $^{+0.1668}_{-0.1855}$ $0.4208$ $_{-0.2555}^{+0.2378}$ $^{+0.2246}_{-0.2077}$
$\mathcal{R}e(A_{f_{0}(500)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $0.6749$ $\pm 0.4192$ $\pm 0.2929$ $1.3057$ $\pm 0.3611$ $\pm 0.2671$
$\mathcal{I}m(A_{f_{0}(500)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ 0.3747-0.3747 $\pm 0.4989$ $\pm 0.2675$ 0.1689-0.1689 $\pm 0.5528$ $\pm 0.2425$
$\mathcal{R}e(A_{f_{0}(980)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ 0.2698-0.2698 $_{-0.2642}^{+0.0870}$ $^{+0.1513}_{-0.1596}$ 0.3532-0.3532 $_{-0.2573}^{+0.0924}$ $^{+0.0725}_{-0.0790}$
$\mathcal{I}m(A_{f_{0}(980)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $1.3567$ $_{-0.2222}^{+0.6411}$ $^{+0.3352}_{-0.3304}$ $0.9319$ $_{-0.1349}^{+0.4909}$ $^{+0.3050}_{-0.3044}$
$\mathcal{R}e(A_{f_{0}(1370)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $1.3511$ $_{-0.5138}^{+0.5718}$ $^{+0.5271}_{-0.5279}$ $0.2864$ $_{-0.5300}^{+0.5436}$ $^{+0.4094}_{-0.4008}$
$\mathcal{I}m(A_{f_{0}(1370)({{K}^{0}_{\mathrm{S}}}\pi)_{\text{S}}})$ $0.1008$ $_{-0.4300}^{+1.0180}$ $^{+0.6857}_{-0.6563}$ $0.5882$ $_{-0.4542}^{+0.4970}$ $^{+0.3376}_{-0.3491}$
$\mathcal{R}e(A_{(a_{1}(1260)\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{0})$ 2.5802-2.5802 $\pm 1.3487$ $\pm 2.2671$ 2.0785-2.0785 $\pm 0.6787$ $\pm 0.9635$
$\mathcal{I}m(A_{(a_{1}(1260)\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{0})$ 0.5676-0.5676 $\pm 0.9027$ $\pm 1.1999$ 0.8797-0.8797 $\pm 0.6419$ $\pm 0.8860$
$\mathcal{R}e(A_{(a_{1}(1260)\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{\parallel})$ 0.1507-0.1507 $\pm 1.5943$ $\pm 0.8448$
$\mathcal{I}m(A_{(a_{1}(1260)\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{\parallel})$ $3.0275$ $\pm 1.3924$ $\pm 1.4278$
$\mathcal{R}e(A_{(a_{1}(1260)\rightarrow f_{0}(500)\pi){{K}^{0}_{\mathrm{S}}}})$ 8.5584-8.5584 $_{-1.8408}^{+1.8330}$ $^{+4.2256}_{-3.9197}$
$\mathcal{I}m(A_{(a_{1}(1260)\rightarrow f_{0}(500)\pi){{K}^{0}_{\mathrm{S}}}})$ $3.5230$ $_{-0.7814}^{+1.2376}$ $^{+2.0371}_{-2.0247}$
$\mathcal{R}e(A_{(a_{1}(1260)\rightarrow f_{0}(980)\pi){{K}^{0}_{\mathrm{S}}}})$ $1.2407$ $_{-1.2627}^{+0.8240}$ $^{+1.3965}_{-1.3885}$
$\mathcal{I}m(A_{(a_{1}(1260)\rightarrow f_{0}(980)\pi){{K}^{0}_{\mathrm{S}}}})$ $0.7200$ $_{-1.0314}^{+0.8666}$ $^{+0.4065}_{-0.4259}$
$\mathcal{R}e(A_{(a_{1}(1640)\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{0})$ 0.3555-0.3555 $\pm 0.4472$ $\pm 0.3954$ 0.2564-0.2564 $\pm 0.4432$ $\pm 0.3951$
$\mathcal{I}m(A_{(a_{1}(1640)\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{0})$ 0.0363-0.0363 $\pm 0.5007$ $\pm 0.3456$ 0.0859-0.0859 $\pm 0.3758$ $\pm 0.2064$
$\mathcal{R}e(A_{(a_{1}(1640)\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{\parallel})$ 0.3892-0.3892 $\pm 0.8504$ $\pm 0.5404$
$\mathcal{I}m(A_{(a_{1}(1640)\rightarrow\rho\pi){{K}^{0}_{\mathrm{S}}}}^{\parallel})$ 0.2174-0.2174 $\pm 0.9290$ $\pm 0.9068$
$\mathcal{R}e(A_{(a_{1}(1640)\rightarrow f_{0}(500)\pi){{K}^{0}_{\mathrm{S}}}})$ $3.8584$ $\pm 1.3145$ $\pm 1.4161$
$\mathcal{I}m(A_{(a_{1}(1640)\rightarrow f_{0}(500)\pi){{K}^{0}_{\mathrm{S}}}})$ $0.5558$ $\pm 1.1537$ $\pm 1.6115$
$\mathcal{R}e(A_{(a_{1}(1640)\rightarrow f_{0}(980)\pi){{K}^{0}_{\mathrm{S}}}})$ 0.3569-0.3569 $_{-0.7193}^{+0.6560}$ $^{+1.1383}_{-1.1493}$
$\mathcal{I}m(A_{(a_{1}(1640)\rightarrow f_{0}(980)\pi){{K}^{0}_{\mathrm{S}}}})$ 1.7988-1.7988 $_{-0.6458}^{+0.6287}$ $^{+1.1608}_{-1.1620}$